Solve the following equations. 
x/3 + 1 = 7/15 
Solve the following equations. 
x/3 + 1 = 7/15 

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

x/3+1=7/15
x/3=7/15–1
x/3 = (7 -15)/15
x/3 = -8/15
x = -8/15 × 3
x=-8/5

Thanks
Shraddha Singh
2 years ago
Dear Student,
We can solve it as follows:
 
X/3 + 1 = 7/15
X/3 = 7/15 - 1
X/3 = (7-15)/15
X/3 = -8/15
15 X = -24
X = -24/15 
X = -8/5
 
Hope it helps!
Thank you ✌️